Jack and Ben cover the same distance at the speeds of 60 km/h and 54 km/h. Find the distance covered by each one of them when Jack takes 30 minutes longer than Ben.

Answer by richwmiller(17219)

As your problem is written, it would be
60(b+.5)=54b
60b+30=54b
6b=-30
b=-5
which is absurd
So it must be that Jack is the slower one at 54 km/h
60b=54(b+.5)
60b=54b+27
6b=27
b=27/6
b=4.5 hr
j=5 hr
60*4.5=270 km for Ben
54*5=270 km for Jack